Solve the following system by solving one of the equations for x or y and then using the substitution method.
x-5y=-12
x-3y=0

Answer by stanbon(75887)

Solve the following system by solving one of the equations for x or y and then using the substitution method.
x-5y=-12
x-3y=0
---------------
x-5y=-12
x = 3y
------
Substitute for "x" and solve for "y":
3y - 5y = -12
-2y = -12
y = 6
----
Solve for "x":
x = 3y
x = 3*6
x = 18
=============
Cheers,
Stan H.